1/ What is the atomic mass and atomic number for the following (show how to write it and express it). How many protons, neutrons, and electrons are present in atoms with the following characteristics?

a. Fe? ?

b. Be? ?

c. Au? ?

d. Pb? ?

2/ How many valence electrons do atoms with the following electron configurations have and what is the chemical symbol?

1s22s22p63s2

1s22s22p63s23p64s23d104p6

1s22s22p63s23p5

1s22s22p63s23p64s23d10
Answer #1

1)

a) Fe

atomic number of Iron(Fe) = 26

atomic mass of Fe = 56

number of protons is always is equal to atomic number of an atom

number of electrtons is equal to atomic number of an atom when it is in neutral

number of neutrons = mass number - atomic number

number of protons = 26

number of neutrons = 56-26 = 30

number of electrons= 26

b) Be

atomic number of Beryllium = 4

atomic mass of Beryllium = 9

number of protons = 4

number of neutrons = 9 - 4 = 5

number of electrons = 4

c) Au

atomic number of Aurum = 79

atomic mass = 197

number of protons = 79

number of neutrons = 197 - 79 = 118

number of electrons = 79

d) Pb

atomic number of Lead = 82

atomic mass of Lead = 207

number of protons = 82

number of neutrons = 207 - 82 = 125

number of electrons = 82

2)

a) 1s22s22p63s2

number of valence electrons = 2 ( 3s2)

Total number of electrons = 12

The symbol of an element = Mg

b) 1s22s22p63s23p64s23d104p6

number of valence electrons = 8 ( 4S2 4P6)

total number of electrons = 36

The symbol of an element = Kr

c) 1s22s22p63s23p5

Number of valence electrons = 7 ( 3s2 3p5)

Total number of electrons = 17

The symbol of an element = Cl

d) 1s22s22p63s23p64s23d10

number of valence electrons = 2 ( 4s2)

total number of valenc electrons = 30

The symbol of an element = Zn